OpenStudy (anonymous):

Solve x + 3y = 9 3x – 3y = –13

jimthompson5910 (jim_thompson5910):

Hint: Add the two equations to get x + 3y = 9 3x – 3y = –13 ------------- 4x + 0y = -4 So 4x = -4 x = ??